# Jet transport coefficients by elastic and radiative scatterings in the strongly interacting quark-gluon plasma

### v1

@@ III Methodology
The propagation of a fast jet parton through the thermalized medium can be characterized by jet transport coefficients that can be calculated within kinetic transport theory by accounting for the sequence of elastic and inelastic interactions of the ”jet” parton per unit length $x$ (or time using $x=ct$, where $c$ is the speed of light. We use the units $c=1$ throughout this work). We note that we do not consider in this study a suppression of the radiation of soft gluons in finite medium due to the LPM effects.
@@
where $p_{i}$ is the 4-momentum of the incoming medium parton, $p_{1}$ and $p_{2}$ are the outgoing jet and medium parton 4-momenta, respectively, $d_{i}$ is the medium parton’s degeneracy factor for spin and color ($2N_{c}$ for quarks and $2(N_{c}^{2}-1)$ for gluons); $f_{i}=f_{i}(E_{i},T,\mu_{q})$ are the Fermi distribution functions for quarks and $f_{i}=f_{i}(E_{i},T)$ are the Bose distribution functions for gluons.
@@
Here the matrix elements squared, i.e. $|\bar{M}_{2\to 2}|^{2}_{\text{jet}+i}$ for elastic $2\to 2$ parton scattering and $|\bar{M}_{2\to 3}|^{2}_{\text{jet}+i}$ for inelastic $2\to 3$ reactions of gluon radiation from the off-shell quark-quark ($q+q$) and quark-gluon ($q+g$) scatterings, are calculated explicitly within the DQPM based on leading order Feynman diagrams with the effective propagators and vertices from the DQPM without any further approximations. The corresponding Feynman diagrams for the $t$-channel of the $q+q\to q+q+g$ and $q+g\to q+g+g$ processes are illustrated in Figure 1. For the details of the evaluation of the $2\to 2$ transport coefficients and for the inelastic $2\to 3$ cross sections we refer the reader to Refs. Grishmanovskii et al. (2022, 2023).
@@
  • $\mathcal{O}=E-E^{\prime}$ – energy loss $\Delta E=dE/dx$ per unit length,
@@
Here $E,E^{\prime}$, $p_{T},p_{T}^{\prime}$ and $p_{L},p_{L}^{\prime}$ denote initial and final energy, transverse and longitudinal momenta of the jet parton, respectively.
Figure 1: Feynman diagrams for the $t$-channel of the $q+q\to q+q+g$ (left) and $q+g\to q+g+g$ (right) processes. The green dots indicate the vertices corresponding to thermal partons. The blue dots indicate the vertices corresponding to the jet quark while the red dots denote the vertices corresponding to the emitted gluon. For the case of the 4-gluon point interaction (diagram 6 in the right figure) the medium parton vertex and the emitted gluon vertex are merged.
@@ III.1 Mass of the radiated gluon
An important feature of the DQPM - contrary to the pQCD-based models - is the non-zero mass of the emitted gluon in the $2\to 3$ reactions. In a thermal DQPM medium the mass of the radiated gluon in the on-shell case is assumed to be the gluon pole mass (obtained from Eq. (4)). The jet parton, however, is not part of the QGP medium and should have a non-thermal mass. In turn, it is not clear whether the emitted gluon mass should still be thermal since the gluon can be emitted from a non-thermal jet parton. Therefore, in this work we also investigate the dependence of transport coefficients on the emitted gluon mass.
@@ III.2 Effective coupling constants for jet elastic and inelastic scattering
In this section we consider five scenarios for different strong couplings for every possible vertex. As illustrated in Fig. 1, for every possible interaction channel for the $2\to 3$ reaction in the $t$-channel of the $q+q\to q+q+g$ (left) and $q+g\to q+g+g$ (right) processes, one can define 3 different strong couplings: one associated with the thermal parton (green dot), one associated with the jet parton (blue dot) and one associated with the emitted gluon (red dot). We note that for the elastic $2\to 2$ reactions the thermal (green) and jet (blue) vertices are the same in the corresponding Feynman diagrams as for the inelastic case.
@@ III.2.1 Scenario 0: DQPM thermal $g(T)$
This scenario corresponds to the default version of the DQPM where in all vertices the thermal DQPM coupling constant is taken, i.e.
@@
Since the DQPM coupling depends strongly on the medium temperature $T$ (cf. blue line in Fig. 2) and grows rapidly in the vicinity of the critical temperature $T_{c}$, we expect a large value and strong $T$-dependence for the temperature dependence of the transport coefficients for inelastic reactions similar to the elastic one, as shown in our previous study Grishmanovskii et al. (2022). The DQPM coupling does not depend on the jet momentum or momentum transfer, so we expect a strong dependence for the inelastic transport coefficients on the initial jet energy $E$ similar to elastic case Grishmanovskii et al. (2022).
@@ III.2.2 Scenario I: $g=const$
As one can see in Fig. 2, although this value of $g$ (gray dotted lines) is smaller than the thermal $g^{DQPM}(T)$ for temperatures up to $T\approx 0.6$ GeV, it is still larger than 1, so it must give a significant contribution to inelastic amplitudes where it is accounted in each vertex. For this scenario we expect a smaller value of $\hat{q}$ compared to ”Scenario 0”, as well as a different form of the temperature dependence and energy dependence of transport coefficients.
@@ III.2.3 Scenario II: $g(Q^{2})$ from the Zakharov model
with $\Lambda_{\text{QCD}}=0.2$ GeV, $Q_{fr}=\Lambda_{\text{QCD}}\exp(2\pi/9\alpha_{s}^{fr})$ while $\alpha_{s}^{fr}$ is a free parameter. In this work we consider $\alpha_{s}^{fr}=1.05$ (vacuum value) and $\alpha_{s}^{fr}=0.42$ (in medium value) taken from Ref. Zakharov (2021). It is important to note that here thermal effects encompassed in $Q^{2}$-dependence suppress the in-medium QCD coupling.
@@
    thermal vertex – $g^{DQPM}(T)$.
@@ III.2.4 Scenario III: QLBT model
Another way to define strong couplings has been introduced in the QLBT model Liu et al. (2021, 2024), where the couplings associated with the jet parton and the emitted gluon vertices are assumed to have the following parametric form:
@@
where $E$ is the jet parton energy in the rest frame, (check!: $T_{c}=150$ MeV) is the critical temperature, and the parameters $A,B$ are determined from the heavy quark observables (such as $R_{AA}$ and $v_{2}$). The coupling for the thermal vertex remains to be $g^{\text{DQPM}}(T)$. The coupling $g^{2}(E)$ is shown in Fig. 2 by the green lines: dashed-dotted for $E=10$ GeV and dashed for $E=100$ GeV on the upped plot showing the temperature dependence and by the dash-dotted line in the lower plot showing the $E$-dependence.
@@
    thermal vertex – $g^{DQPM}(T)$.
@@ III.2.5 Scenario IV: DREENA framework
The last scenario is motivated by the couplings used in the DREENA framework Zigic et al. (2022); Karmakar et al. (2023) where the jet interaction with the thermal QCD medium is computed within the two loop HTL model Djordjevic and Djordjevic (2014); Djordjevic (2009) and implies the following form of the coupling:
@@
    thermal vertex – $g^{DQPM}(T)$.
@@
The orange lines in Fig. 2 show the coupling $g^{2}(E)$: dot-dashed for $E=10$ GeV and dashed for $E=100$ GeV on the upped plot for the $T$-dependence and by the dot-dashed line for $T=0.2$ GeV and by the dashed line for $T=0.4$ GeV in the lower plot for the $E$-dependence.
@@ IV Results: radiative differential cross sections
Figure 3: DQPM differential cross sections $d\sigma/dk_{t}$ as a function of the transverse momentum $k_{t}$ of the emitted gluon (upper left), $d\sigma/dq_{t}$ as a function of the transverse momentum $q_{t}$ of the jet (upper right), $d\sigma/dE_{g}$ as a function of the energy $E_{g}$ of the emitted gluon (lower left), and $x\ d\sigma/dx$ as a function of the longitudinal ratio $x=p_{L}^{g}/p_{L}^{j}$ (lower right) for $q+q\to q+q+g$ scattering at $\sqrt{s}=10$ GeV and $T=0.3$ GeV in case of massive (thermal) emitted gluon (blue lines) and massless emitted gluon (orange lines). The green line on the upper left plot shows $d\sigma/dk_{t}$ calculated with the pQCD BAMPS model for $\alpha_{s}=0.3$ and massless partons without LPM cutoff.
We recall that the $2\to 3$ cross sections and interaction rates for $q+q\to q+q+g$ and $q+g\to q+g+g$ scattering have been calculated within the DQPM in Ref. Grishmanovskii et al. (2023) and compared to the elastic one. It has been shown that the inelastic reactions contribute only little to the interaction rate $\Gamma$ of the QGP in thermal equilibrium. However, their contribution is expected to be important for the calculation of the energy loss of the off-equilibrium fast jet parton propagating through the thermal QGP medium, since the inelastic $q+q$ and $q+g$ cross sections grow with increasing collision energy Grishmanovskii et al. (2023).
@@
In Fig. 3 we present the different types of differential DQPM cross sections for the case of thermal massive – with the pole DQPM mass $M_{g}(T)$ defined by Eq. (4) – (blue lines) and massless (orange lines) emitted gluon for the $q+q\to q+q+g$ scattering at $\sqrt{s}=10$ GeV and $T=0.3$ GeV. The upper left plot shows the differential cross sections $d\sigma/dk_{t}$ as a function of the transverse momentum $k_{t}$ of the emitted gluon. The upper right plot indicates the differential cross section $d\sigma/dq_{t}$ as a function of the transverse momentum $q_{t}$ of the jet (upper right). The lower left plot displays the differential cross section $d\sigma/dE_{g}$ as a function of the energy $E_{g}$ of the emitted gluon. The lower left plot demonstrates the distribution $x\ d\sigma/dx$ as a function of the longitudinal ratio $x=p_{L}^{g}/p_{L}^{j}$, where $p_{L}^{g}$ and $p_{L}^{j}$ are the longitudinal momenta of gluon and initial jet parton, respectively.
One can see from Fig. 3 all DQPM differential cross sections for radiation of massless gluons at small $k_{t},q_{t},E_{g}$ or $x$ are larger than for the massive cases. This is due to the fact that the kinematically available phase space is strongly reduced for the emission of massive heavy gluons compared to the massless ones. Therefore, it is expected to obtain a similar increase of transport coefficients for massless emitted gluons versus massive ones. Our result is consistent with the findings in Refs. Gossiaux et al. (2010); Aichelin et al. (2014); Gossiaux et al. (2014) for the radiative energy loss of heavy quarks in the QGP medium described by scalar pQCD, where it has been shown that the gluon emission cross section decreases with increasing gluon mass.
In the upper left plot we compare the DQPM $d\sigma/dk_{t}$ to the pQCD differential cross section within the BAMPS framework for $\alpha_{s}=0.3$ and massless partons (green line) which diverges at $k_{t}\to 0$ without LPM effect Senzel (2020). Thus, the pQCD $d\sigma/dk_{t}$ is much larger at low $k_{t}$ and decreases much faster with increasing $k_{t}$ relative to the DQPM cross sections.
@@ V Results: $\hat{q}$ and energy loss $dE/dx$
Here we present the $T$- and $p$-dependence of the jet transport coefficients – $\hat{q}$ and energy loss $dE/dx$ – for elastic and inelastic reactions for different scenarios for the strong coupling as presented in Section III.2 and compare our results with previous estimations from the literature. Also we study the dependence of transport coefficients on the mass of emitted gluons within the DQPM.
We will show the results for the ’elastic’ and ’inelastic’ (or radiative) $\hat{q}$ as well as for the sum of both elastic and inelastic contributions:

### v2

@@ III Methodology
The propagation of a fast jet parton through the thermalized medium can be characterized by jet transport coefficients that can be calculated within kinetic transport theory by accounting for the sequence of elastic and inelastic interactions of the jet parton per unit length $x$ (or time, using $x=ct$, where $c$ is the speed of light; we use the units $c=1$ throughout this work). We note that in this study we do not consider a suppression of the radiation of soft gluons in finite medium due to the LPM effects.
@@
where $p_{i}$ is the 4-momentum of the incoming medium parton, $p_{1}$ and $p_{2}$ are the outgoing jet and medium parton 4-momenta, respectively; $d_{i}$ is the medium parton’s degeneracy factor for spin and color ($2N_{c}$ for quarks and $2(N_{c}^{2}-1)$ for gluons); $f_{i}=f_{i}(E_{i},T,\mu_{q})$ are the Fermi distribution functions for quarks, and $f_{i}=f_{i}(E_{i},T)$ are the Bose distribution functions for gluons.
@@
  • $\mathcal{O}=E-E^{\prime}$ – energy loss $dE/dx$ per unit length,
@@
Here $E,E^{\prime}$, $p_{T},p_{T}^{\prime}$, and $p_{L},p_{L}^{\prime}$ denote initial and final energy, transverse and longitudinal momenta of the jet parton, respectively.
In Eq. (9) the matrix elements squared, i.e., $|\bar{M}_{2\to 2}|^{2}_{\text{jet}+i}$ for elastic $2\to 2$ parton scattering and $|\bar{M}_{2\to 3}|^{2}_{\text{jet}+i}$ for inelastic $2\to 3$ reactions with gluon radiation from the off-shell quark-quark ($q+q$) and quark-gluon ($q+g$) scatterings, are calculated explicitly within the DQPM based on leading order Feynman diagrams with the effective propagators and vertices from the DQPM without any further approximations.
The corresponding Feynman diagrams for the $t$ channel of the $q+q\to q+q+g$ and $q+g\to q+g+g$ processes are illustrated in Fig. 1. Considering that the jet parton is a $u$-quark, the following inelastic reactions and channels are accounted for in this study:
  • $u+u\to u+u+g$ ($t+u$ channels)
  • $u+\bar{u}\to u+\bar{u}+g$ ($t+s$ channels)
  • $u+d\to u+d+g$ ($t$ channel)
  • $u+\bar{d}\to u+\bar{d}+g$ ($t$ channel)
  • $u+s\to u+s+g$ ($t$ channel)
  • $u+\bar{s}\to u+\bar{s}+g$ ($t$ channel)
  • $u+\bar{u}\to d+\bar{d}+g$ ($s$ channel)
  • $u+\bar{u}\to s+\bar{s}+g$ ($s$ channel)
  • $u+g\to u+g+g$ ($t+u+s$ channels).
We note that the $u$ channel for the $u+u\to u+u+g$ reaction gives practically identical contribution as the $t$ channel and is accounted for by doubling the $t$ channel contribution when evaluating the transport coefficient to simplify the calculations. The $s$ channels for all above reactions are strongly suppressed relative to $t$ channels and give only minor contributions to the transport coefficients. In case of the $u+g\to u+g+g$ reaction the $u$ channel is also strongly suppressed (cf. Ref. Grishmanovskii et al. (2024)). Therefore, only the $t$ channels are considered in the current study. For the details of the evaluation of the $2\to 2$ transport coefficients and for the inelastic $2\to 3$ cross sections we refer the reader to Refs. Grishmanovskii et al. (2022, 2024).
Figure 1: Feynman diagrams for the $t$ channel of the $q+q\to q+q+g$ (left) and $q+g\to q+g+g$ (right) processes. The green dots indicate the vertices corresponding to thermal partons. The blue dots indicate the vertices corresponding to the jet quark while the red dots denote the vertices corresponding to the emitted gluon. For the case of the 4-gluon point interaction (diagram 6 in the right figure) the medium parton vertex and the emitted gluon vertex are merged.
@@ III.1 Mass of the radiated gluon
An important feature of the DQPM, contrary to the pQCD-based models, is the non-zero mass of the emitted gluon in the $2\to 3$ reactions. In a thermal DQPM medium the mass of the radiated gluon in the on-shell case is assumed to be the gluon pole mass (obtained from Eq. (4)). The jet parton, however, is not part of the QGP medium and should have a non-thermal mass. In turn, it is not clear whether the emitted gluon mass should still be thermal since the gluon can be emitted from a non-thermal jet parton. Therefore, in this work we also investigate the dependence of transport coefficients on the emitted gluon mass.
@@ III.2 Effective coupling constants for jet elastic and inelastic scattering
In this section, we consider five ”scenarios” for different strong couplings for every possible vertex. As illustrated in Fig. 1, for every possible interaction channel for the $2\to 3$ reaction in the $t$ channel of the $q+q\to q+q+g$ (left) and $q+g\to q+g+g$ (right) processes, one can define three different strong couplings: one associated with the thermal parton (green dot), one associated with the jet parton (blue dot), and one associated with the emitted gluon (red dot). We note that for the elastic $2\to 2$ reactions the thermal (green) and jet (blue) vertices are the same in the corresponding Feynman diagrams as for the inelastic case.
@@ III.2.1 Scenario 0: DQPM thermal $g(T)$
This scenario corresponds to the default version of the DQPM where in all vertices the thermal DQPM coupling constant is taken, i.e.,
@@
Since the DQPM coupling depends strongly on the medium temperature $T$ (cf. blue line in Fig. 2) and grows rapidly in the vicinity of the critical temperature $T_{c}$, we expect a large value and strong temperature dependence of the transport coefficients for inelastic reactions similar to the elastic one, as shown in our previous study Grishmanovskii et al. (2022). The DQPM coupling does not depend on the jet momentum or momentum transfer, so we expect a strong dependence for the inelastic transport coefficients on the initial jet energy $E$ similar to elastic case Grishmanovskii et al. (2022).
@@ III.2.2 Scenario I: $g=const$
As one can see in Fig. 2, although this value of $g$ (gray dotted lines) is smaller than the thermal $g^{\mathrm{DQPM}}(T)$ for temperatures up to $T\approx 0.6$ GeV, it is still larger than 1, so it must give a significant contribution to inelastic amplitudes, where it is accounted in each vertex. For this scenario, we expect a smaller value of $\hat{q}$ compared to Scenario 0, as well as a different form of the temperature dependence and energy dependence of transport coefficients.
@@ III.2.3 Scenario II: $g(Q^{2})$ from the Zakharov model
with $\Lambda_{\text{QCD}}=0.2$ GeV, $Q_{fr}=\Lambda_{\text{QCD}}\exp(2\pi/9\alpha_{s}^{fr})$, and $\alpha_{s}^{fr}$ is a free parameter. In this work, we consider $\alpha_{s}^{fr}=1.05$ (vacuum value) and $\alpha_{s}^{fr}=0.42$ (in-medium value) taken from Ref. Zakharov (2021). It is important to note that here thermal effects encompassed in $Q^{2}$ dependence suppress the in-medium QCD coupling.
@@
    thermal vertex – $g^{\mathrm{DQPM}}(T)$.
@@ III.2.4 Scenario III: QLBT model
Another way to define strong couplings has been introduced in the QLBT model Liu et al. (2022, 2024), where the couplings associated with the jet parton and the emitted gluon vertices are assumed to have the following parametric form:
@@
where $E$ is the jet parton energy in the rest frame, $T_{c}=150$ MeV is the critical temperature, and the parameters $A,B$ are determined from the heavy quark observables (such as $R_{AA}$ and $v_{2}$). The coupling for the thermal vertex remains to be $g^{\text{DQPM}}(T)$. The coupling $g^{2}(E)$ is shown in Fig. 2 by the green lines: dashed-dotted for $E=10$ GeV and dashed for $E=100$ GeV on the upper plot, showing the temperature dependence, and by the dash-dotted line in the lower plot, showing the jet energy dependence.
@@
    thermal vertex – $g^{\mathrm{DQPM}}(T)$.
@@ III.2.5 Scenario IV: DREENA framework
The last scenario is motivated by the couplings used in the DREENA framework Zigic et al. (2022); Karmakar et al. (2023), where the jet interaction with the thermal QCD medium is computed within the two-loop HTL model Djordjevic and Djordjevic (2014); Djordjevic (2009) and implies the following form of the coupling:
@@
    thermal vertex – $g^{\mathrm{DQPM}}(T)$.
@@
The orange lines in Fig. 2 show the coupling $g^{2}(E)$: dot-dashed for $E=10$ GeV and dashed for $E=100$ GeV on the upped plot for the $T$ dependence, and dot-dashed for $T=0.2$ GeV and dashed for $T=0.4$ GeV in the lower plot for the $E$ dependence.
@@ IV Results: radiative differential cross sections
Figure 3: DQPM differential cross sections: $d\sigma/dk_{t}$ as a function of the transverse momentum $k_{t}$ of the emitted gluon (upper left), $d\sigma/dq_{t}$ as a function of the transverse momentum $q_{t}$ of the jet (upper right), $d\sigma/dE_{g}$ as a function of the energy $E_{g}$ of the emitted gluon (lower left), and $x\ d\sigma/dx$ as a function of the longitudinal ratio $x=p_{L}^{g}/p_{L}^{j}$ (lower right) for $q+q\to q+q+g$ scattering at $\sqrt{s}=10$ GeV and $T=0.3$ GeV in case of massive (thermal) emitted gluon (blue lines) and massless emitted gluon (orange lines). The green line on the upper left plot shows $d\sigma/dk_{t}$ calculated with the pQCD BAMPS model for $\alpha_{s}=0.3$ and massless partons without LPM cutoff.
We recall that the $2\to 3$ cross sections and interaction rates for $q+q\to q+q+g$ and $q+g\to q+g+g$ scattering have been calculated within the DQPM in Ref. Grishmanovskii et al. (2024) and compared to the elastic ones. It has been shown that the inelastic reactions contribute only little to the interaction rate $\Gamma$ of the QGP in thermal equilibrium. That is due to the fact that partonic scatterings at larger $\sqrt{s}$ are strongly suppressed in the thermal QGP – cf. Fig. 17 of Ref. Grishmanovskii et al. (2024), thus, only low-$\sqrt{s}$ elastic and inelastic reactions dominantly contribute to the interaction rates, where the elastic cross sections are much larger than the inelastic ones. However, inelastic contribution is expected to be important for the calculation of the energy loss of the off-equilibrium fast jet parton propagating through the thermal QGP medium, since the inelastic $q+q$ and $q+g$ cross sections grow with increasing collision energy Grishmanovskii et al. (2024).
@@
In Fig. 3 we present the different types of differential DQPM cross sections for the case of thermal massive – with the pole DQPM mass $M_{g}(T)$ defined by Eq. (4) – (blue lines) and massless (orange lines) emitted gluon for the $q+q\to q+q+g$ scattering at $\sqrt{s}=10$ GeV and $T=0.3$ GeV. The upper left plot shows the differential cross sections $d\sigma/dk_{t}$ as a function of the transverse momentum $k_{t}$ of the emitted gluon. The upper right plot indicates the differential cross section $d\sigma/dq_{t}$ as a function of the transverse momentum $q_{t}$ of the jet. The lower left plot displays the differential cross section $d\sigma/dE_{g}$ as a function of the energy $E_{g}$ of the emitted gluon. The lower right plot demonstrates the distribution $x\ d\sigma/dx$ as a function of the longitudinal ratio $x=p_{L}^{g}/p_{L}^{j}$, where $p_{L}^{g}$ and $p_{L}^{j}$ are the longitudinal momenta of gluon and initial jet parton, respectively.
One can see from Fig. 3, all DQPM differential cross sections for radiation of massless gluons at small $k_{t},q_{t},E_{g}$, or $x$ are larger than for the massive cases. This is due to the fact that the kinematically available phase space is strongly reduced for the emission of massive heavy gluons compared to the massless ones. Therefore, it is expected to obtain a similar increase of transport coefficients for massless emitted gluons versus massive ones. Our results are consistent with the findings in Refs. Gossiaux et al. (2010); Aichelin et al. (2014); Gossiaux et al. (2014) for the radiative energy loss of heavy quarks in the QGP medium described by scalar pQCD, where it has been shown that the gluon emission cross section decreases with increasing gluon mass.
In the upper left plot we compare the DQPM $d\sigma/dk_{t}$ to the pQCD differential cross section within the BAMPS framework with $\alpha_{s}=0.3$ and massless partons (green line), which diverges at $k_{t}\to 0$ without LPM effect Senzel (2020). Thus, the pQCD $d\sigma/dk_{t}$ is much larger at low $k_{t}$ and decreases much faster with increasing $k_{t}$ relative to the DQPM cross sections.
@@ V Results: $\hat{q}$ and energy loss $dE/dx$
Here we present the $T$ and $p$ dependence of the jet transport coefficients – $\hat{q}$ and energy loss $dE/dx$ – for elastic and inelastic reactions for different scenarios for the strong coupling as presented in Sec. III.2 and compare our results with previous estimations from the literature. Also we study the dependence of transport coefficients on the mass of emitted gluons within the DQPM.
We will show the results for the elastic and inelastic (or radiative) $\hat{q}$ coefficient as well as for the sum of both elastic and inelastic contributions: